# OMNI-AVSR: Towards Unified Multimodal Speech Recognition with Large Language Models

### v1

@@ I Introduction
Auditory Speech Recognition (ASR) [amodei2016deep, watanabe2017hybrid, prabhavalkar2023end] often degrades in noisy environments such as crowded areas or subways. To address this limitation, Audio-Visual Speech Recognition (AVSR) [dupont2000audio, potamianos2004audio, petridis2018end] incorporates visual cues, such as lip movements, which remain unaffected by acoustic noise, thereby enhancing recognition robustness and accuracy. Early AVSR methods relied on modality-specific encoders and handcrafted fusion strategies [noda2015audio, mroueh2015deep, petridis2018audio]. The introduction of Transformers [vaswani2017attention] significantly advanced performance [afouras2018deep, ma2021end, serdyuk2021audio], spurring research into multimodal learning paradigms such as self-supervision [shi2022learning, shi2022robust, haliassos2023jointly], ASR-to-AVSR distillation [rouditchenko2024whisper, autoavsr], and cross-modal complementarity [hong2023watch, chen2023leveraging].
More recently, Multimodal Large Language Models (MLLMs) have demonstrated that integrating modalities such as vision and speech significantly extends the capabilities of LLMs, yielding state-of-the-art results across diverse tasks [bai2025qwen2, fini2025multimodal, yao2024dense, fathullah2024prompting, goel2025audio, wu2025step]. Building on this progress, several studies have applied LLMs to ASR, Visual Speech Recognition (VSR), and AVSR, with promising results [llama-AVSR, llama-SMoP, yeo2025mms, yeo2025zero, yeo2024visual, llama-MTSK].
However, most existing approaches treat each task in isolation, training separate models for ASR, VSR, and AVSR. This not only increases computational cost and complexity but also overlooks potential synergies across tasks. In contrast, studies across multiple domains have demonstrated the feasibility of unified multi-task multimodal LLMs [zhu2024uni, unifiedmllm, wu2024omni, xu2025qwen2, zhang2025stream]. While some attempts have been made to unify ASR, VSR, and AVSR, these either rely on cumbersome student-teacher pseudo-labeling frameworks [haliassos2024unified] or underperform compared to task-specific models [hsu2022u, torrie2025multiavsr].
Motivated by these limitations, we introduce Omni-AVSR, a unified audio-visual LLM capable of performing ASR, VSR, and AVSR within a single framework. To adapt the backbone LLM to all tasks in a parameter-efficient manner, we propose three LoRA-based methods. Furthermore, we adapt and optimize the matryoshka representation learning paradigm [kusupati2022matryoshka, cai2024matryoshka, llama-MTSK] for our setting, enabling efficient multi-granularity training while mitigating its inherent computational cost. This allows the number of tokens to be dynamically adjusted at inference according to resource availability and task requirements. To the best of our knowledge, Omni-AVSR is the first audio-visual LLM that supports ASR, VSR, and AVSR jointly while enabling elastic inference under a single set of weights.
@@ II Omni-AVSR
Following prior audio-visual LLMs [llama-AVSR, llama-MTSK, llama-SMoP, yeo2025mms], Omni-AVSR comprises pre-trained audio and video encoders, projection layers, and an LLM backbone (see Fig. 1a). In the next sections, we detail how Omni-AVSR is endowed with 1) explicit control over audio-visual granularities during inference and 2) the ability to jointly support ASR, VSR, and AVSR within a single model.
@@ II-A Multi-Granularity via Efficient Matryoshka Training
Given an audio waveform $\mathbf{a}$ and its corresponding lip movement video $\mathbf{v}$, we process them with a pre-trained audio encoder (e.g., Whisper [radford2023robust]) and video encoder (e.g., AV-HuBERT [shi2022learning]) to obtain audio and visual tokens, $\mathbf{Z}^{a}$ and $\mathbf{Z}^{v}$, respectively. Reducing token granularity lowers computational cost and improves efficiency when feeding audio-visual tokens into an LLM. In AVSR, temporal continuity across modalities creates redundancy, yet most compression methods rely on fixed rates, limiting adaptability to performance and resource trade-offs [llama-AVSR, llama-SMoP, yeo2025mms]. While finer-grained tokens enhance recognition accuracy, they substantially increase inference compute cost due to the quadratic complexity of Transformers.
To address this, Llama-MTSK [llama-MTSK] exploits the matryoshka representation learning (MRL) principle [cai2024matryoshka] to flexibly control audio-visual granularity at inference based on user requirements. During training, token sequences at varying granularities are generated by applying $C_{A}$ audio compression rates $a_{1},a_{2},\cdots,a_{C_{A}}$} and $C_{V}$ video rates $\{v_{1},v_{2},\cdots,v_{C_{V}}$ to the input streams. For AVSR, each of the $C_{A}\cdot C_{V}$ audio-visual sequences is fed to the LLM, requiring $C_{A}\cdot C_{V}$ forward/backward passes per batch.
@@ II-C Efficient LLM Adaptation via Omni-LoRA
In Omni-AVSR, following prior works [llama-AVSR, llama-SMoP, llama-MTSK, yeo2024visual, yeo2025mms], the pre-trained LLM is kept frozen while low-rank LoRA modules [hu2022lora] are employed to parameter-efficiently fine-tune it. Given our multi-task setting, we explore three configurations: 1) Omni-LoRA-S, 2) Omni-LoRA-T, and 3) Omni-LoRA-ST, illustrated in Fig. 1b. These variants allow us to systematically investigate the trade-off between parameter sharing and task specialization within Omni-AVSR.
@@ III Experiments and Results
Llama-AVSR [llama-AVSR] 3.3 4.3 26.9 30.0 2.5 2.6 3.9 4.6 9.8
Llama-MTSK [llama-MTSK] 2.5 3.9 26.7 28.5 2.5 2.5 3.7 4.0 9.3
@@
Llama-AVSR [llama-AVSR] 1.1 2.0 27.4 29.5 1.1 1.2 2.0 2.1 8.3
Llama-MTSK [llama-MTSK] 1.0 2.0 26.9 27.8 1.0 1.0 1.9 2.0 8.0
@@
Llama-AVSR [llama-AVSR] $C_{A}$ + $C_{V}$ + $C_{A}C_{V}$ $C_{A}$ + $C_{V}$ + $C_{A}C_{V}$
Llama-MTSK [llama-MTSK] $T$ $C_{A}$ + $C_{V}$ + $C_{A}C_{V}$
@@ III-A Experiment Settings
Datasets. We conduct experiments on LRS2 [chung2017lip] and LRS3 [afouras2018lrs3] datasets. LRS2 includes $225$ hours of footage from BBC programs. LRS3 contains $433$ hours of English video clips from TED talks.
Pre-Processing. We follow [autoavsr, llama-AVSR, llama-MTSK] for the datasets pre-processing. For video, we crop the mouth region of interests (ROIs) through a bounding box of $96$ × $96$. Each frame is normalised by subtracting the mean and dividing by the standard deviation of the training set. Audio data undergo z-normalisation per utterance.
Omni-AVSR Details. We use AV-HuBERT Large as the visual encoder and Whisper medium as the audio encoder. The projection layers consist of two linear layers with a ReLU activation in between. For the LLM backbone, we adopt LLaMA 3.2-1B [dubey2024llama] in our main experiments. Following prior work [llama-AVSR, llama-MTSK, llama-SMoP], both the LLM and video encoder are fine-tuned via LoRA modules applied to the query and value projection matrices with rank $64$. We evaluate three Omni-AVSR variants, depending on the LoRA configuration used: Omni-AVSR-S, Omni-AVSR-T, and Omni-AVSR-ST.
Training/Inference Details. Following [autoavsr, llama-AVSR, llama-MTSK], we augment visual inputs through horizontal flipping, random cropping, and adaptive time masking, while for audio we only apply adaptive time masking. We define the textual prompts as in [llama-AVSR, llama-MTSK, llama-SMoP]: “Transcribe {task_prompt} to text.”, where task_prompt $\in$ {“speech”, “video”, “speech and video”}. We set $\lambda_{\mathsf{ASR}}=\lambda_{\mathsf{AVSR}}=1$ and $\lambda_{\mathsf{VSR}}=1.5$. We train our Omni-AVSR models for $8$ epochs with the AdamW optimizer with cosine annealing scheduler and weight decay set to $0.1$ using NVIDIA L40 GPUs. The learning rate is 1e-3. For decoding, we use beam search with a beam width of $15$ and temperature of $0.6$.
Audio-Visual Granularities. For fair comparison with prior work [llama-MTSK], we adopt the same compression rates, chosen to capture a spectrum of efficiency–performance trade-offs at inference. Specifically, we use {4,16} for ASR, {2,5} for VSR, and their Cartesian product for AVSR, yielding four audio-visual configurations. Token compression is performed via average pooling [llama-MTSK].
Baselines. As shown in Tables I and III, we compare Omni-AVSR variants with three main approaches: 1) Llama-AVSR [llama-AVSR], which trains a separate model for each task and compression rate; 2) Llama-MTSK [llama-MTSK], which enables elastic inference by training on multiple rates but only within a single task; 3) Llama-MT, which supports multi-task (MT) training across ASR, VSR, and AVSR but requires a separate model for each rate. In contrast, Omni-AVSR unifies both elastic inference and multi-task learning within a single framework, subsuming these baselines as special cases. Additional comparisons with AVSR sota methods are provided in Section III-C.
@@
Llama-AVSR [llama-AVSR] 2.6 4.1 4.8 12.1 19.1
Llama-MTSK [llama-MTSK] 2.5 3.9 4.8 11.7 18.5
@@
Llama-AVSR [llama-AVSR] 4.2 5.8 6.5 14.9 22.1
Llama-MTSK [llama-MTSK] 3.8 5.5 6.0 14.0 20.5
@@
Table IV: Comparison with state-of-the-art methods using a single model for ASR, VSR, and AVSR on LRS3. u-HuBERT is trained on LRS3 and VoxCeleb2, totaling $1759$ hours.
@@
u-HuBERT [hsu2022u] 325 1759 1.5 29.1 1.3
MultiAVSR [torrie2025multiavsr] 274 433 2.4 31.1 2.5
USR [haliassos2024unified] 171 433 1.9 34.3 1.6
@@ III-B Main Results
Table I reports the ASR/VSR/AVSR results of our three Omni-AVSR variants on LRS2 and LRS3. On LRS2, the task-specific variant Omni-AVSR-T achieves the best performance, while on LRS3 all three variants yield comparable results. This difference is likely due to the larger training set of LRS3, which enables lower WERs overall, particularly for ASR and AVSR. Compared with the baselines, we observe the following: (1) all Omni-AVSR variants consistently outperform Llama-AVSR, which requires a separate model per rate and task; (2) Omni-AVSR-T on LRS2, and all three variants on LRS3, match or surpass Llama-MTSK and Llama-MT, with Omni-AVSR-S and -T attaining average WERs as low as $7.9$ across tasks on LRS3; (3) task-wise, Omni-AVSR particularly benefits VSR; and (4) performance trends remain consistent across compression rates. These results demonstrate that Omni-AVSR delivers competitive or superior accuracy while unifying elastic inference and multi-task learning within a single framework.
@@
Results under Acoustic Noise. To evaluate the robustness of Omni-AVSR under noisy conditions, we inject babble noise from the NOISEX dataset [noisex] at varying SNRs. As shown in Table III, Omni-AVSR-ST consistently outperforms Llama-AVSR and Llama-MTSK, and remains competitive with Llama-MT across noise levels, often surpassing it at lower SNRs.
@@
In Table IV, we compare Omni-AVSR-ST with three state-of-the-art methods that train a single model for ASR, VSR, and AVSR: u-HuBERT [hsu2022u], MultiAVSR [torrie2025multiavsr], and USR [haliassos2024unified]. Unlike Omni-AVSR, these methods do not support elastic inference. At the (4,2) compression setting, Omni-AVSR-ST achieves the best performance across all tasks while requiring significantly fewer parameters and surpassing u-HuBERT, despite the latter being trained on $1759$ hours of data (LRS3 + VoxCeleb2 datasets). Even under the more extreme (16,5) compression, Omni-AVSR-ST maintains competitive results within a single set of weights.
@@ III-C Ablation Studies
AVSR Comparison with Sota Methods. Fig. 2 (left) presents a comparison of Omni-AVSR-ST with recent state-of-the-art approaches on LRS3 for the AVSR task. Baselines include UniVPM [hu2023hearing], USR [haliassos2024unified], Whisper-Flamingo [rouditchenko2024whisper], Llama-AVSR [llama-AVSR], Auto-AVSR [autoavsr], AV-HuBERT [shi2022learning], and MMS-Llama [yeo2025mms]. Omni-AVSR-ST (evaluated at audio-video rates of (4,2)) achieves competitive WERs while requiring substantially fewer parameters and training data hours than all baselines, within one consistent framework.
LLM Scaling Trend. We study how scaling the LLM impacts performance in Fig. 2 (right). Specifically, we evaluate Llama 3.2–1B, 3B, and 3.1–8B [dubey2024llama], as well as Qwen 2.5–0.5B, 1.5B, 3B, 7B, 14B, and 32B [qwen2.5]. Results are reported for ASR at audio rate 16 (black outline), VSR at video rate 2 (violet outline), and AVSR at (4,2) rates. As shown, performance improves with larger LLMs, with higher gains observed on more challenging tasks (e.g., VSR) or under higher compression (e.g., ASR at rate 16). However, larger models incur greater training computations, memory usage, and slower inference. Overall, LLMs in the 1–3B parameter range represent a favorable trade-off between accuracy and efficiency.

### v2

@@ I Introduction
Auditory Speech Recognition (ASR) [3, 40, 30] often degrades in noisy environments such as crowded areas or subways. To address this limitation, Audio-Visual Speech Recognition (AVSR) [12, 29, 28] incorporates visual cues, such as lip movements, which remain unaffected by acoustic noise, thereby enhancing recognition robustness and accuracy. Early AVSR methods relied on modality-specific encoders and handcrafted fusion strategies [26, 25, 27]. The introduction of Transformers [39] significantly advanced performance [1, 23, 34], spurring research into multimodal learning paradigms such as self-supervision [35, 36, 15], ASR-to-AVSR distillation [33, 24], and cross-modal complementarity [17, 9].
More recently, Multimodal Large Language Models (MLLMs) have demonstrated that integrating modalities such as vision and speech significantly extends the capabilities of LLMs, yielding state-of-the-art results across diverse tasks [4, 45, 13, 14, 41]. Building on this progress, several studies have applied LLMs to ASR, Visual Speech Recognition (VSR), and AVSR, with promising results [7, 8, 47, 48, 46, 6].
However, most existing approaches treat each task in isolation, training separate models for ASR, VSR, and AVSR. This not only increases computational cost and complexity but also overlooks potential synergies across tasks. In contrast, studies across multiple domains have demonstrated the feasibility of unified multi-task multimodal LLMs [50, 22, 42, 43, 49]. While some attempts have been made to unify ASR, VSR, and AVSR, these either rely on cumbersome student-teacher pseudo-labeling frameworks [16] or underperform compared to task-specific models [18, 32, 37].
Motivated by these limitations, we introduce Omni-AVSR, a unified audio-visual LLM capable of performing ASR, VSR, and AVSR within a single framework. To adapt the backbone LLM to all tasks in a parameter-efficient manner, we propose three LoRA-based methods. Furthermore, we adapt and optimize the matryoshka representation learning paradigm [21, 5, 6] for our setting, enabling efficient multi-granularity training while mitigating its inherent computational cost. This allows the number of tokens to be dynamically adjusted at inference according to resource availability and task requirements. To the best of our knowledge, Omni-AVSR is the first audio-visual LLM that supports ASR, VSR, and AVSR jointly while enabling elastic inference under a single set of weights.
@@ II Omni-AVSR
Following prior audio-visual LLMs [7, 6, 8, 47], Omni-AVSR comprises pre-trained audio and video encoders, projection layers, and an LLM backbone (see Fig. 1a). In the next sections, we detail how Omni-AVSR is endowed with 1) explicit control over audio-visual granularities during inference and 2) the ability to jointly support ASR, VSR, and AVSR within a single model.
@@ II-A Multi-Granularity via Efficient Matryoshka Training
Given an audio waveform $\mathbf{a}$ and its corresponding lip movement video $\mathbf{v}$, we process them with a pre-trained audio encoder (e.g., Whisper [31]) and video encoder (e.g., AV-HuBERT [35]) to obtain audio and visual tokens, $\mathbf{Z}^{a}$ and $\mathbf{Z}^{v}$, respectively. Reducing token granularity lowers computational cost and improves efficiency when feeding audio-visual tokens into an LLM. In AVSR, temporal continuity across modalities creates redundancy, yet most compression methods rely on fixed rates, limiting adaptability to performance and resource trade-offs [7, 8, 47]. While finer-grained tokens enhance recognition accuracy, they substantially increase inference compute cost due to the quadratic complexity of Transformers.
To address this, Llama-MTSK [6] exploits the matryoshka representation learning (MRL) principle [5] to flexibly control audio-visual granularity at inference based on user requirements. During training, token sequences at varying granularities are generated by applying $C_{A}$ audio compression rates $a_{1},a_{2},\cdots,a_{C_{A}}$} and $C_{V}$ video rates $\{v_{1},v_{2},\cdots,v_{C_{V}}$ to the input streams. For AVSR, each of the $C_{A}\cdot C_{V}$ audio-visual sequences is fed to the LLM, requiring $C_{A}\cdot C_{V}$ forward/backward passes per batch.
@@ II-C Efficient LLM Adaptation via Omni-LoRA
In Omni-AVSR, following prior works [7, 8, 6, 46, 47], the pre-trained LLM is kept frozen while low-rank LoRA modules [19] are employed to parameter-efficiently fine-tune it. Given our multi-task setting, we explore three configurations: 1) Omni-LoRA-S, 2) Omni-LoRA-T, and 3) Omni-LoRA-ST, illustrated in Fig. 1b. These variants allow us to systematically investigate the trade-off between parameter sharing and task specialization within Omni-AVSR.
@@ III Experiments and Results
Llama-AVSR [7] 3.3 4.3 26.9 30.0 2.5 2.6 3.9 4.6 9.8
Llama-MTSK [6] 2.5 3.9 26.7 28.5 2.5 2.5 3.7 4.0 9.3
@@
Llama-AVSR [7] 1.1 2.0 27.4 29.5 1.1 1.2 2.0 2.1 8.3
Llama-MTSK [6] 1.0 2.0 26.9 27.8 1.0 1.0 1.9 2.0 8.0
@@
Llama-AVSR [7] $C_{A}$ + $C_{V}$ + $C_{A}C_{V}$ $C_{A}$ + $C_{V}$ + $C_{A}C_{V}$
Llama-MTSK [6] $T$ $C_{A}$ + $C_{V}$ + $C_{A}C_{V}$
@@ III-A Experiment Settings
Datasets. We conduct experiments on LRS2 [10] and LRS3 [2] datasets. LRS2 includes $225$ hours of footage from BBC programs. LRS3 contains $433$ hours of English video clips from TED talks.
Pre-Processing. We follow [24, 7, 6] for the datasets pre-processing. For video, we crop the mouth region of interests (ROIs) through a bounding box of $96$ × $96$. Each frame is normalised by subtracting the mean and dividing by the standard deviation of the training set. Audio data undergo z-normalisation per utterance.
Omni-AVSR Details. We use AV-HuBERT Large as the visual encoder and Whisper medium as the audio encoder. The projection layers consist of two linear layers with a ReLU activation in between. For the LLM backbone, we adopt LLaMA 3.2-1B [11] in our main experiments. Following prior work [7, 6, 8], both the LLM and video encoder are fine-tuned via LoRA modules applied to the query and value projection matrices with rank $64$.
Training/Inference Details. Following [24, 7, 6], we augment visual inputs through horizontal flipping, random cropping, and adaptive time masking, while for audio we apply adaptive time masking. We set the textual prompts as in [7, 6, 8]: “Transcribe {task_prompt} to text.”, where task_prompt $\in$ {“speech”, “video”, “speech and video”}. We set $\lambda_{\mathsf{ASR}}=\lambda_{\mathsf{AVSR}}=1$ and $\lambda_{\mathsf{VSR}}=1.5$. We train our Omni-AVSR models for $8$ epochs with AdamW optimizer with cosine annealing scheduler and weight decay set to $0.1$. The learning rate is 1e-3. For decoding, we use beam search with a beam width of $15$ and temperature of $0.6$.
Audio-Visual Granularities. For fair comparison with prior work [6], we adopt the same compression rates, chosen to capture a spectrum of efficiency–performance trade-offs at inference. Specifically, we use {4,16} for ASR, {2,5} for VSR, and their Cartesian product for AVSR, yielding four audio-visual configurations. Token compression is performed via average pooling [6].
Baselines. As shown in Tables I and III, we compare Omni-AVSR variants with three main approaches: 1) Llama-AVSR [7], which trains a separate model for each task and compression rate; 2) Llama-MTSK [6], which enables elastic inference by training on multiple rates but only within a single task; 3) Llama-MT, which supports multi-task (MT) training across ASR, VSR, and AVSR but requires a separate model for each rate. In contrast, Omni-AVSR unifies both elastic inference and multi-task learning within a single framework, subsuming these baselines as special cases. Additional comparisons with AVSR sota methods are provided in Section III-C.
@@
Llama-AVSR [7] 2.6 4.1 4.8 12.1 19.1
Llama-MTSK [6] 2.5 3.9 4.8 11.7 18.5
@@
Llama-AVSR [7] 4.2 5.8 6.5 14.9 22.1
Llama-MTSK [6] 3.8 5.5 6.0 14.0 20.5
@@
Table IV: Comparison with state-of-the-art methods using a single model for ASR, VSR, and AVSR on LRS3. u-HuBERT and AV-CPL are trained on LRS3 and VoxCeleb2, totaling $1759$ hours.
@@
u-HuBERT [18] 325 1759 1.5 29.1 1.3
AV-CPL [32] 325 1759 2.3 47.4 2.2
MultiAVSR [37] 274 433 2.4 31.1 2.5
USR [16] 171 433 1.9 34.3 1.6
@@ III-B Main Results
Table I reports the ASR/VSR/AVSR results of our three Omni-AVSR variants on LRS2 and LRS3. On LRS2, among the three proposed variants, Omni-AVSR-T achieves the best performance, while on LRS3 all three variants yield comparable results. This difference is likely due to the larger training set of LRS3, which enables lower WERs overall, particularly for ASR and AVSR. Compared with the baselines, we observe the following: (1) all Omni-AVSR variants consistently outperform Llama-AVSR, which requires a separate model per rate and task; (2) Omni-AVSR-T on LRS2, and all three variants on LRS3, match or surpass Llama-MTSK and Llama-MT, with Omni-AVSR-S and -T attaining average WERs as low as $7.9$ across tasks on LRS3; (3) task-wise, Omni-AVSR particularly benefits VSR; and (4) performance trends remain consistent across compression rates. These results demonstrate that Omni-AVSR delivers competitive or superior accuracy while unifying elastic inference and multi-task learning within a single framework.
@@
Results under Acoustic Noise. To evaluate the robustness of Omni-AVSR under noisy conditions, we inject babble noise from the NOISEX dataset [38] at varying SNRs. As shown in Table III, Omni-AVSR-ST consistently outperforms Llama-AVSR and Llama-MTSK, and remains competitive with Llama-MT across noise levels, often surpassing it at lower SNRs.
@@
In Table IV, we compare Omni-AVSR-ST with three state-of-the-art methods that train a single model for ASR, VSR, and AVSR: u-HuBERT [18], AV-CPL , MultiAVSR [37], and USR [16]. Unlike Omni-AVSR, these methods do not support elastic inference. At the (4,2) compression setting, Omni-AVSR-ST achieves the best performance across all tasks while requiring significantly fewer parameters and surpassing u-HuBERT and AV-CPL, despite both being trained on $1759$ hours of data (LRS3 + VoxCeleb2 datasets). Even under the more extreme (16,5) compression, Omni-AVSR-ST maintains competitive results within a single set of weights.
@@ III-C Ablation Studies
AVSR Comparison with Sota Methods. Fig. 2 (left) presents a comparison of Omni-AVSR-ST with recent state-of-the-art approaches on LRS3 for the AVSR task. Baselines include UniVPM [20], USR [16], Whisper-Flamingo [33], Llama-AVSR [7], Auto-AVSR [24], AV-HuBERT [35], and MMS-Llama [47]. Omni-AVSR-ST (evaluated at audio-video rates of (4,2)) achieves competitive WERs while requiring substantially fewer parameters and training data hours than all baselines, within one consistent framework.
LLM Scaling Trend. We study how scaling the LLM impacts performance in Fig. 2 (right). Specifically, we evaluate Llama 3.2–1B, 3B, and 3.1–8B [11], as well as Qwen 2.5–0.5B, 1.5B, 3B, 7B, 14B, and 32B [44]. Results are reported for ASR at audio rate 16 (black outline), VSR at video rate 2 (violet outline), and AVSR at (4,2) rates. As shown, performance improves with larger LLMs, with higher gains observed on more challenging tasks (e.g., VSR) or under higher compression (e.g., ASR at rate 16). However, larger models incur greater training computations, memory usage, and slower inference. Overall, LLMs in the 1–3B parameter range represent a favorable trade-off between accuracy and efficiency.